5. (16) A which the sample simulated exercise gave n 23 observations on escape time (sec) for oil workers, from mean and sample standard deviation are 372.99 and 22.51, respectively ippose the investigators had believed a priori that true average escape time would be at most 6 m Does the data contradict this prior belief? Assuming normality, test the appropriate hypotheses using a significa nce level of 0.0S. a) State the appropriate hypotheses. b) Calculate the test statistic and determine the P-value. (Round your test statistic to two places and your P-value to three decimal places.) decimal
Answer #1

5)

Solution :

This is the right tailed test .

The null and alternative hypothesis is ,

H0 :  \mu\leq 6

Ha : \mu > 6

\bar x = 372.99 / 60 = 6.2165

\mu = 6

s = 0.375167

n = 23

df = n - 1 = 23 - 1 = 22

Test statistic = t = (\bar x - \mu) / s / \sqrtn

= (6.2165 - 6) /0.375167 / \sqrt23 = 2.77

Test statistic = 2.77

This is the right tailed test .

P-value = 0.0056

\alpha = 0.05

P-value < \alpha

Reject the null hypothesis .

Cliam is true .